OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. x^2 − 16x + ____

OpenStudy (imstuck):

This is another way of asking you to complete the square. Do you know how?

OpenStudy (anonymous):

Not a clue.

OpenStudy (anonymous):

A perfect square trinomial has one factor that is squared. For example: \[(x+4)^{2}\rightarrow(x-4)^{2}\rightarrow (x+1)\] And so on. Each of those is a perfect square trinomial because it can be written with the same factor twice (x+4)(x+4) (x-4)(x-4) (x+1)(x+1) All of that being said, we are looking for what factor we can multiply by the same factor again to get the beginning part to be \[x ^{2}-16x...\]

OpenStudy (imstuck):

Take half the x term and square it. -16/2 = -8. -8 squared is 64. So add 64 in and then you have a perfect square binomial, like this:\[x ^{2}-16x+64 = (x-8)^{2}\]

OpenStudy (anonymous):

The (x+1) example up at the top is supposed to be squared like \[(x+1)^{2}\]

OpenStudy (imstuck):

The coefficient in front of the x^2 term HAS TO BE A 1, then you will always take half the x term and square it. THat's how to complete the square.

OpenStudy (anonymous):

I'm really confused

OpenStudy (imstuck):

about what?

OpenStudy (anonymous):

So, we ask ourselves, what factor can be added together with itself to give us -16. That would be -8. So we know (x-8)(x-8) would give us \[x ^{2}-16x... \] Then what @BassCatcher15 ?

OpenStudy (anonymous):

These are all just examples of perfect square trinomials... \[(x+1)^{2}\rightarrow(x-5)^{2}\rightarrow(x+13)^{2}\rightarrow(x+4)^{2}\]

OpenStudy (nikato):

complete the square normally you see trinomials written like this \[ax ^{2}+bx+c\] but right now you are only given \[ax ^{2}+bx\] and you're trying to find c

OpenStudy (anonymous):

64?

OpenStudy (anonymous):

You got it.

OpenStudy (nikato):

to find c, divide b by 2 and square it so \[(\frac{ b }{ 2 })^{2}=c\]

OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. (Do not type the variable in the blank.) 4x^2 + ___x + 49 This would be 7^2 right?

OpenStudy (anonymous):

@nikato

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (nikato):

7^2 is not right

OpenStudy (anonymous):

Then can you explain how to do this?

OpenStudy (nikato):

can you think of a number that equals 49 when its squared?

OpenStudy (anonymous):

7

OpenStudy (nikato):

yes. and the same with 4x^2. what squared equals 4x^2

OpenStudy (anonymous):

jeez louise what is half of sixteen?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

yes what is eight squared?

OpenStudy (anonymous):

64

OpenStudy (anonymous):

yes that is your answer done finished stick a fork in it

OpenStudy (anonymous):

Haha. Thanks!

OpenStudy (anonymous):

yw

OpenStudy (nikato):

i thought we were done with that question already

OpenStudy (anonymous):

I thought so too. Did the next question not post?

OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. (Do not type the variable in the blank.) 4x^2 + ___x + 49

OpenStudy (nikato):

yea. i was trying to help you. so what number squared equals 4x^2

OpenStudy (anonymous):

16 right?

OpenStudy (nikato):

no. 2x because (2x)^2 = 4x^2 16=4^2

OpenStudy (anonymous):

Ohhhh okay!

OpenStudy (nikato):

so if we were to factor this PERFECT trinomial it would be \[(2x ^{2}+7)^{2}\]

OpenStudy (mathstudent55):

\(4x^2 + \_\_\_\_x + 49\) The \(4x^2\) term comes form \(2x \times 2x\) The \(49\) term comes from \(7 \times 7\) Then the trinomial must come from \((2x + 7)(2x + 7) \) Let's square \((2x + 7):\) \( (2x + 7)^2 = 4x^2 + 28x + 49\) Now you see clearly what the x term is.

OpenStudy (anonymous):

So the blank is 7?

OpenStudy (mathstudent55):

Look at my last line. What's the coefficient of the x term? It's 28.

OpenStudy (anonymous):

Oh okay. Thanks!

OpenStudy (mathstudent55):

\(4x^2 + \_\_\_\_x + 49\) \(4x^2 + \color{red}{28}x + 49\)